# How Performance Pressure Influences AI-Assisted Decision Making

### v2

@@ 5.1 AI advice reliance under pressure
The RAIR decrease in LIME in the baseline pressure condition compared to no-XAI suggests LIME decreased trust in the AI advice. It associated individual words with a prediction that may not have made sense (e.g., “is” is associated with being genuine; 3(b)), but users reported associating typos, excessive punctuation, and grammar with their prediction (§4.1). Payment pressure appears to mitigate the negative effects of LIME, improving RAIR over the LIME baseline (but not over the no-XAI baseline) without changing RSR. This result points toward the potential for XAI techniques in the same style as LIME to help mitigate or further improve AI-assisted decision making behavior on tasks where humans are close to random.
@@
One XAI method can be more effective than another depending on the task or stakeholders Jiang et al. (2022), and our results show that this is the case even under different pressure conditions. Furthermore, the negative effect of pressure can override the potential benefits of XAI, as seen in the GenAI RAIR decrease.
Time pressure has a largely consistent, negative impact on RAIR, RSR, and overall performance. The behavior change mirrors the phenomenon of “choking,” where a human experiences performance decline in critical high pressure settings, often attributed to anxiety. Participants were anxious about completing the tasks within the 30 second time limit, and rushed to complete them faster (averaging only 20 seconds, compared to 30 when working without time pressure). As a result, they relied heavily on the AI advice. One subject stated: “For the timed cases, I ended up relying on the AI’s decision and just went with it. However, for the other parts, I looked out for typos and missing punctuations to detect human reviews.”
Swaroop et al. (2024) observed that slowing down AI overreliers in higher time pressure environments could improve their performance, and since our participants had plenty of time to spare, we expect such a strategy would be helpful in this setting. Emotional regulation strategies can be effective in improving performance when anxiety is high Balk et al. (2013), which can help alleviate anxiety over running out of time (and hence the possibility of earning a bonus). Methods from the distraction model may also help: if choking occurs due to cognitive overload (rather than anxiety), an attention shift could refocus the user to pay attention to relevant information and avoid choking Hardy et al. (2001); Mullen et al. (2005); Eysenck (2012); Nieuwenhuys and Oudejans (2012). For example, adding a third color to LIME explanations, unrelated to the prediction, could be a distractor.
@@ 6 Future Directions
Our work has illustrated how environmental stressors and XAI can alter AI-assisted decision making behavior. The mixed effectiveness of incorporating XAI under pressures reinforces the findings of Jiang et al. (2022) and Swaroop et al. (2024) where AI advice or XAI are only effective for some people, some environments, and at certain times. This calls for building adaptive environments that can be personalized for the user and task to provide appropriate XAI methods, cognitive forcing functions, and other AI advice interventions. Such environments could improve AI-assisted decision making and help elicit more diverse data for training robust NLP models. However, in order to develop these environments, we need a deeper understanding of how environmental factors interact with user characteristics and different XAI methods. Future work should explore other pressures such as competition and multi-tasking, alongside XAI and measures of user personality and behavior, e.g., how they cope with stress. Additional work investigating the influence of performance pressure on tasks where humans have higher than random chance performance should also be investigated, as the literature suggests performance pressure can improve performance, despite our results suggesting minimal influence.
@@ 7 Conclusion
AI assistance is already used in the real world for tasks with high and low risk, and designing adaptive AI assistants that are domain-specific requires understanding the different factors influencing how humans use AI advice. We investigated how pressure influences the use of AI advice. Using deceptive review classification, crowdworkers with little expertise or personal motivation requirements, and two different XAI techniques, we observed complex effects on how pressure influences AI advice usage. Pressure and XAI interactions could both improve and decrease appropriate AI advice usage along the two dimensions of appropriate AI reliance and appropriate self-reliance. While performance pressure had minor effects, time pressure had a strong negative effect. Our results contribute to the body of literature investigating how pressure influences AI-assisted decision making. We note the relevance of these findings in AI annotation projects in particular; our work motivates continued research on the effects of pressure on AI assistance in varied environments while taking into account individual differences.
@@ A.1 RAIR and RSR
    $CAIR$: 1 if the initial judgment disagrees with the ground truth, the AI advice is correct, and the final judgment agrees with the ground truth; 0 otherwise
@@
    $CA$: 1 if the initial judgment is incorrect and the AI advice is correct, regardless of the final judgment; 0 otherwise
@@
    $CSR$: 1 if the initial judgment agrees is correct, the AI advice is incorrect, and the final judgment is correct; 0 otherwise
@@
    $IA$: 1 if the initial judgment is correct and the AI advice is incorrect; 0 otherwise
@@ A.2 Presurvey
    MPQ
  • I would never go hang gliding or bungee jumping.
  • I would never make a high-risk investment.
  • I avoid dangerous situations.
  • I seek danger.
  • I am willing to try anything once.
  • I do dangerous things.
  • I enjoy being reckless
  • I seek adventure
@@ A.3.1 Holt-Laury Survey
Appendix B Regression Tables
The random effects model regression tables are detailed below.
DV: RAIR Coef. Std.Err. z P>|z| [0.025 0.975]
Intercept 0.548 0.059 9.310 0.000 0.432 0.663
Pressure -0.022 0.012 -1.761 0.078 -0.046 0.002
Explanation Condition -0.008 0.024 -0.334 0.738 -0.056 0.040
MPQ -0.075 0.041 -1.812 0.070 -0.156 0.006
Holt-Laury -0.059 0.043 -1.367 0.172 -0.143 0.025
Advisor Trust -0.193 0.061 -3.158 0.002** -0.312 -0.073
Trust in AI -0.016 0.018 -0.923 0.356 -0.051 0.018
Age -0.023 0.020 -1.141 0.254 -0.062 0.016
Race 0.003 0.009 0.380 0.704 -0.014 0.020
Gender 0.040 0.034 1.177 0.239 -0.027 0.107
Education 0.014 0.007 1.949 0.051 -0.000 0.029
Subject Group 0.083 0.038
Table 2: DV: RAIR. * $p<0.1$, ** $p<0.05$, ***$p<0.01$
DV: RSR Coef. Std.Err. z P>|z| [0.025 0.975]
Intercept 0.618 0.062 10.017 0.000 0.497 0.739
Pressure -0.043 0.012 -3.595 $0$0.066 -0.020
Explanation Condition -0.069 0.026 -2.686 0.007** -0.119 -0.019
MPQ 0.051 0.044 1.161 0.246 -0.035 0.136
Holt-Laury 0.044 0.045 0.974 0.330 -0.045 0.133
Advisor Trust 0.164 0.064 2.559 0.011* 0.038 0.290
Trust in AI 0.005 0.018 0.294 0.768 -0.031 0.042
Age 0.014 0.021 0.666 0.506 -0.027 0.055
Race -0.004 0.009 -0.471 0.638 -0.022 0.013
Gender -0.035 0.036 -0.986 0.324 -0.106 0.035
Education -0.005 0.008 -0.589 0.556 -0.020 0.011
Subject Group 0.098 0.044
Table 3: DV: RSR. * $p<0.1$, ** $p<0.05$, ***$p<0.01$
DV: Overall Accuracy Coef. Std.Err. z P>|z| [0.025 0.975]
Intercept 0.598 0.018 33.481 0.000 0.563 0.632
Pressure -0.016 0.006 -2.721 0.007** -0.028 -0.005
Explanation Condition -0.036 0.007 -5.096 $0$0.050 -0.022
MPQ 0.000 0.012 0.040 0.968 -0.023 0.024
Holt-Laury -0.006 0.013 -0.489 0.625 -0.031 0.018
Advisor Trust 0.005 0.018 0.260 0.795 -0.030 0.040
Trust in AI -0.002 0.005 -0.383 0.701 -0.012 0.008
Age -0.006 0.006 -1.036 0.300 -0.018 0.005
Race -0.002 0.003 -0.774 0.439 -0.007 0.003
Gender -0.006 0.010 -0.622 0.534 -0.026 0.013
Education 0.005 0.002 2.362 0.018* 0.001 0.009
Subject Group 0.003 0.007
Table 4: DV: Accuracy. * $p<0.1$, ** $p<0.05$, ***$p<0.01$

### v3

@@ 5.1 AI advice reliance under pressure
The RAIR decrease in LIME in the baseline pressure condition compared to no-XAI suggests LIME decreased trust in the AI advice.666 Disentangling AI trust and AI reliance behavior is its own challenge (Scharowski et al., 2022; Guo et al., 2024) It associated individual words with a prediction that may not have made sense (e.g., “is” is associated with being genuine; 3(b)), but users reported associating typos, excessive punctuation, and grammar with their prediction (§4.1). Payment pressure appears to mitigate the negative effects of LIME, improving RAIR over the LIME baseline (but not over the no-XAI baseline) without changing RSR. This result points toward the potential for XAI techniques in the same style as LIME to help mitigate or further improve AI-assisted decision making behavior on tasks where humans are close to random.
@@
One XAI method can be more effective than another depending on the task or stakeholders (Jiang et al., 2022), and our results show that this is the case even under different pressure conditions. Furthermore, the negative effect of pressure can override the potential benefits of XAI, as seen in the GenAI RAIR decrease.
Time pressure has a largely consistent, negative impact on RAIR, RSR, and overall performance. The behavior change mirrors the phenomenon of “choking,” where a human experiences performance decline in critical high pressure settings, often attributed to anxiety. Participants were anxious about completing the tasks within the 30 second time limit, and rushed to complete them faster. As a result, they relied heavily on the AI advice. One subject stated: “For the timed cases, I ended up relying on the AI’s decision and just went with it. However, for the other parts, I looked out for typos and missing punctuations to detect human reviews.”
Swaroop et al. (2024) observed that slowing down AI overreliers in higher time pressure environments could improve their performance, and since our participants had plenty of time to spare, we expect such a strategy would be helpful in this setting. Emotional regulation strategies can be effective in improving performance when anxiety is high (Balk et al., 2013), which can help alleviate anxiety over running out of time (and hence the possibility of earning a bonus). Methods from the distraction model may also help: if choking occurs due to cognitive overload (rather than anxiety), an attention shift could refocus the user to pay attention to relevant information and avoid choking (Hardy et al., 2001; Mullen et al., 2005; Eysenck, 2012; Nieuwenhuys and Oudejans, 2012). For example, adding a third color to LIME explanations, unrelated to the prediction, could be a distractor.
@@ 6 Future Directions
Our work has illustrated how environmental stressors and XAI can alter AI-assisted decision making behavior. The mixed effectiveness of incorporating XAI under pressures reinforces the findings of Jiang et al. (2022) and Swaroop et al. (2024) where AI advice or XAI are only effective for some people, some environments, and at certain times. The challenges we observed in determining significant predictive factors of AI advice taking behavior also reflect the difficulty in modeling appropriate AI advice reliance. This calls for building adaptive environments that can be personalized for the user and task to provide appropriate XAI methods, cognitive forcing functions, and other AI advice interventions. Such environments could improve AI-assisted decision making and help elicit more diverse data for training robust NLP models. However, in order to develop these environments, we need a deeper understanding of how environmental factors interact with user characteristics and different XAI methods. Future work should explore other pressures such as competition and multi-tasking, alongside XAI and measures of user personality and behavior, e.g., how they cope with stress. Additional work investigating the influence of performance pressure on tasks where humans have higher than random chance performance should also be investigated, as the literature suggests performance pressure can improve performance, despite our results suggesting minimal influence.
@@ 7 Conclusion
AI assistance is already used in the real world for tasks with high and low risk, and designing adaptive AI assistants that are domain-specific requires understanding the different factors influencing how humans use AI advice. We investigated how pressure influences the use of AI advice. Using deceptive review classification, crowdworkers with little expertise or personal motivation requirements, and two different XAI techniques, we observed complex effects on how pressure influences AI advice usage. Pressure and XAI interactions could both improve and decrease appropriate AI advice. While performance pressure had minor effects, time pressure had a strong negative effect. Our results contribute to the body of literature investigating AI-assisted decision making. We note the relevance of these findings in AI annotation projects in particular; our work motivates continued research on the effects of pressure on AI assistance in varied environments while taking into account individual differences.
@@
Acknowledgements
This research used statistical consulting resources provided by the Center for Statistics and the Social Sciences, University of Washington. We thank Dr. Thomas Frye for valuable insights in leveraging methods from behavioral economics.
@@ A.1 RAIR and RSR
    CAIRCAIR: 1 if the initial judgment disagrees with the ground truth, the AI advice is correct, and the final judgment agrees with the ground truth; 0 otherwise
@@
    CACA: 1 if the initial judgment is incorrect and the AI advice is correct, regardless of the final judgment; 0 otherwise
@@
    CSRCSR: 1 if the initial judgment agrees is correct, the AI advice is incorrect, and the final judgment is correct; 0 otherwise
@@
    IAIA: 1 if the initial judgment is correct and the AI advice is incorrect; 0 otherwise
@@ A.2 Presurvey
  1. 1.
    I would never go hang gliding or bungee jumping.
  2. 2.
    I would never make a high-risk investment.
  3. 3.
    I avoid dangerous situations.
  4. 4.
    I seek danger.
  5. 5.
    I am willing to try anything once.
  6. 6.
    I do dangerous things.
  7. 7.
    I enjoy being reckless
  8. 8.
    I seek adventure
@@ A.3.1 Holt-Laury Survey
Appendix B Results
XAI Pressure Pre-Advice Post-Advice
NoXAI Baseline 0.52 0.55
Payment 0.54 0.55
Time 0.51 0.52
Overall 0.52 0.54
LIME Baseline 0.52 0.55
Payment 0.54 0.55
Time 0.51 0.52
Overall 0.52 0.54
GenAI Baseline 0.52 0.55
Payment 0.54 0.55
Time 0.51 0.52
Overall 0.52 0.54
Table 2: Accuracy before and after receiving AI advice. Accuracy marginally increased after receiving AI advice.
B.1 Significance Tests
A Shapiro test for normality indicates our data is not normal. To test for significance, we use the Friedman test with Wilcoxon post-hoc tests for our within-subjects (pressure condition) data and the Kruskal-Wallis and Dunn post-hoc tests for our between-subjects (XAI condition) data. We report uncorrected and Holm-Bonferroni corrected p-values when significant.
Pressure condition tests.
For RAIR, when GenAI was present, the difference between baseline and time pressure tasks had p-values of $p=.009$, corrected to $p=.028$. For RSR, when GenAI was present, there was a significant difference between baseline and payment pressures ($p=.008$, corrected $p=.016$), and baseline and time pressures ($p=.003$, corrected $p=.009$).
XAI condition tests.
For RSR, GenAI significantly differed from both the baseline no-XAI ($p=8\mathrm{e}{-6}$, corrected $p=2.4\mathrm{e}{-5}$) and LIME ($p=3.4\mathrm{e}{-4}$, corrected $p=.0007$). Similarly, for overall task accuracy, GenAI significantly differed from no-XAI ($p=1.7\mathrm{e}{-7}$, corrected $p=5.0\mathrm{e}{-7}$) and LIME ($p=7\mathrm{e}{-7}$, corrected $p=1\mathrm{e}{-6}$).
B.2 Regression Tables
The generalized linear mixed models using a Gaussian distribution and log-transformed dependent variable results are detailed below. Values in parentheses were transformed back to the original scale.
DV: RAIR Mean ($\beta$) SD HDI 3% HDI 97% MCSE Mean MCSE SD ESS Bulk ESS Tail $\hat{R}$
Sigma 2.101 0.050 2.005 2.192 0.000 0.001 11390.000 6135.000 1.000
Intercept -1.523 0.225 -1.943 -1.095 0.002 0.003 12491.000 6218.000 1.000
Pressure -0.107 (0.899) 0.085 -0.269 (0.764) 0.049 (1.05) 0.001 0.001 11686.000 5580.000 1.000
Explanation -0.07 (0.932) 0.088 -0.232 (0.793) 0.097 (1.102) 0.001 0.001 11010.000 6630.000 1.000
MPQ* -0.332 (0.717) 0.151 -0.608 (0.544) -0.047 (0.954) 0.001 0.002 11254.000 6139.000 1.000
Holt-Laury* -0.391 (0.676) 0.155 -0.682 (0.506) -0.095 (0.909) 0.001 0.002 12901.000 6782.000 1.000
Advisor Trust* -0.858 (0.424) 0.218 -1.248 (0.287) -0.432 (0.649) 0.002 0.002 11377.000 6740.000 1.000
Trust in AI -0.066 (0.936) 0.064 -0.188 (0.829) 0.052 (1.053) 0.001 0.001 10956.000 6341.000 1.000
Age -0.121 (0.886) 0.071 -0.249 (0.78) 0.014 (1.014) 0.001 0.001 11665.000 6379.000 1.000
Race 0.015 (1.015) 0.031 -0.042 (0.959) 0.073 (1.076) 0.000 0.000 11200.000 6051.000 1.000
Gender 0.167 (1.182) 0.122 -0.061 (0.941) 0.392 (1.48) 0.001 0.001 11282.000 6049.000 1.000
Education* 0.072 (1.075) 0.027 0.02 (1.02) 0.12 (1.127) 0.000 0.000 11259.000 5936.000 1.000
Table 3: GLMM results for RAIR. * indicates credible effect (97% HDI excludes 1). Mean and HDI values in parentheses have been transformed back to the original scale.
DV: RSR Mean ($\beta$) SD HDI 3% HDI 97% MCSE Mean MCSE SD ESS Bulk ESS Tail $\hat{R}$
Sigma 2.027 0.048 1.936 2.116 0.000 0.001 12155.000 6525.000 1.000
Intercept -1.394 0.214 -1.802 -1.003 0.002 0.003 12251.000 6168.000 1.000
Pressure* -0.169 (0.845) 0.084 -0.331 (0.718) -0.015 (0.985) 0.001 0.001 12549.000 6463.000 1.000
Explanation* -0.24 (0.787) 0.085 -0.397 (0.672) -0.074 (0.929) 0.001 0.001 12539.000 6400.000 1.000
MPQ 0.223 (1.25) 0.143 -0.038 (0.963) 0.491 (1.634) 0.001 0.002 11592.000 6475.000 1.000
Holt-Laury 0.111 (1.117) 0.147 -0.158 (0.854) 0.391 (1.478) 0.001 0.002 12765.000 6687.000 1.000
Advisor Trust* 0.738 (2.092) 0.211 0.33 (1.391) 1.125 (3.08) 0.002 0.002 10998.000 6235.000 1.000
Trust in AI 0.029 (1.029) 0.060 -0.079 (0.924) 0.147 (1.158) 0.001 0.001 11932.000 6689.000 1.000
Age 0.053 (1.054) 0.069 -0.076 (0.927) 0.182 (1.2) 0.001 0.001 11936.000 6174.000 1.000
Race 0.0 (1.0) 0.029 -0.055 (0.946) 0.053 (1.054) 0.000 0.000 12330.000 6358.000 1.000
Gender* -0.23 (0.795) 0.115 -0.447 (0.64) -0.018 (0.982) 0.001 0.001 13571.000 6405.000 1.000
Education -0.019 (0.981) 0.025 -0.067 (0.935) 0.029 (1.029) 0.000 0.000 10931.000 6202.000 1.000
Table 4: GLMM results for RSR. * indicates credible effect (97% HDI excludes 1). Mean and HDI values in parentheses have been transformed back to the original scale.
DV: Overall Accuracy Mean ($\beta$) SD HDI 3% HDI 97% MCSE Mean MCSE SD ESS Bulk ESS Tail $\hat{R}$
Sigma 0.317 0.008 0.303 0.331 0.000 0.000 11965.000 6325.000 1.000
Intercept -0.532 0.034 -0.598 -0.472 0.000 0.000 11254.000 6552.000 1.000
Pressure* -0.032 (0.969) 0.013 -0.056 (0.946) -0.008 (0.992) 0.000 0.000 12719.000 5529.000 1.000
Explanation* -0.06 (0.942) 0.013 -0.084 (0.919) -0.034 (0.967) 0.000 0.000 11629.000 6508.000 1.000
MPQ -0.011 (0.989) 0.023 -0.052 (0.949) 0.033 (1.034) 0.000 0.000 12045.000 6433.000 1.000
Holt-Laury -0.018 (0.982) 0.023 -0.06 (0.942) 0.025 (1.025) 0.000 0.000 11983.000 6352.000 1.000
Advisor Trust -0.013 (0.987) 0.034 -0.077 (0.926) 0.05 (1.051) 0.000 0.000 11390.000 6084.000 1.000
Trust in AI 0.002 (1.002) 0.009 -0.017 (0.983) 0.019 (1.019) 0.000 0.000 11321.000 6503.000 1.000
Age -0.009 (0.991) 0.011 -0.029 (0.971) 0.012 (1.012) 0.000 0.000 11287.000 6351.000 1.000
Race -0.003 (0.997) 0.005 -0.012 (0.988) 0.006 (1.006) 0.000 0.000 12220.000 6107.000 1.000
Gender -0.025 (0.975) 0.018 -0.059 (0.943) 0.011 (1.011) 0.000 0.000 12496.000 6420.000 1.000
Education 0.007 (1.007) 0.004 -0.0 (1.0) 0.015 (1.015) 0.000 0.000 11597.000 6548.000 1.000
Table 5: GLMM results for overall accuracy. * indicates credible effect (97% HDI excludes 1). Mean and HDI values in parentheses have been transformed back to the original scale.